OpenStudy (anonymous):

What transformation has changed the parent function f(x) = 3(2)x to its new appearance shown in the graph below?

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

f(x)+2 f(x)-2 f(x+2) f(x-2) These are the answer choices

jhonyy9 (jhonyy9):

sorry but i not can open this site can you upload it here please

OpenStudy (anonymous):

did u get it now

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

so than you see the graph where the graph interesect the axe y on unit 1 how many will be x for this point the intersection ?

OpenStudy (anonymous):

- .5

jhonyy9 (jhonyy9):

when this diagram intersect the axe y how many will be there x ?

jhonyy9 (jhonyy9):

y there is equal 1

jhonyy9 (jhonyy9):

so do you know the coordinates points (0,0) what mean ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

x=? y=?

OpenStudy (anonymous):

i know y is (0,1)

jhonyy9 (jhonyy9):

yes right

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

i hate graphs

jhonyy9 (jhonyy9):

but than you check now these choices so i think this point (0,1) you not can edit it with them - yes ? try it please

jhonyy9 (jhonyy9):

so than you give to x value zero you need getting for y = 1 and i see there not is possible with indifferent what use it for - yes ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

wait moment please for x=1 result y=4 can you get it anywhere ?

OpenStudy (anonymous):

is it going to be f(x+2)

OpenStudy (anonymous):

there's an intersection point at (1,4)

OpenStudy (anonymous):

and one at (-.5,1)

OpenStudy (anonymous):

another at (-6,-2) i'm guessing

jhonyy9 (jhonyy9):

check it please -5 is with -2 too same -6

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so you said that believ it f(x+2) why ? how you prove it ?

OpenStudy (anonymous):

i just guessed... x=1 and i divide 4 by2

jhonyy9 (jhonyy9):

sorry what mean this 3(2)x this 2 inside parentheses wan mean ?

jhonyy9 (jhonyy9):

what mean - sorry

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

f(x)+2 f(x)-2 f(x+2) f(x-2)

jhonyy9 (jhonyy9):

no for the function f(x) = 3(2)x what is 2 inside parentheses here next 3 ?

OpenStudy (anonymous):

y

jhonyy9 (jhonyy9):

is exponent ?

jhonyy9 (jhonyy9):

or you dont know it really ?

OpenStudy (anonymous):

the x is an exponent

OpenStudy (anonymous):

3(2)^x

jhonyy9 (jhonyy9):

so there are 3*2^x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it going to = 6^x

jhonyy9 (jhonyy9):

no this is wrong because the exponent is just for 2 and for 3 not

jhonyy9 (jhonyy9):

so than there is 3*2^x for x=0 result how many ?

OpenStudy (anonymous):

0?

jhonyy9 (jhonyy9):

no you dont know that every integers on exponent - power - zero equal ?

jhonyy9 (jhonyy9):

equal one

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

anything to 0=1

jhonyy9 (jhonyy9):

so in tis case of your exercise how many will result the exponent zero ?

OpenStudy (anonymous):

1

jhonyy9 (jhonyy9):

there are 3*2^0 = ?

OpenStudy (anonymous):

6

jhonyy9 (jhonyy9):

above you have said that 2^0 = 1

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

3*2^0=3

jhonyy9 (jhonyy9):

yes so than check the graph for what function will get these choices - one from these choices right ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

f(x-2)?

jhonyy9 (jhonyy9):

why ? can you prove it with points from the diagram ?

jhonyy9 (jhonyy9):

for x=1 f(x-2)=f(1-2)=f(-1) = 3(2)^-1 = ?

jhonyy9 (jhonyy9):

not is right for x=1 you need getting that you see this diagram f(1)= 4

OpenStudy (anonymous):

confused!

jhonyy9 (jhonyy9):

sorry - sorry

jhonyy9 (jhonyy9):

i see it now but you need prove it how you have got f(x-2) right sure - because you have wrote it above - can you prove it now for me here ?

OpenStudy (anonymous):

can u show me how to prove it

OpenStudy (anonymous):

im confused on graphs

OpenStudy (anonymous):

is f(x-2) correct?

jhonyy9 (jhonyy9):

f(x)=3*2^x f(0)=3*1=3 f(x-2) = 3*2^x = f(x) not is - sorry

jhonyy9 (jhonyy9):

f(x)+2 f(x)-2 f(x+2) f(x-2) f(x)= 3*2^x f(1)= 3*2=6 sorry i not get it right - you have wrote your exercise right sure and these choices too ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's ok i'll just guess an answer

OpenStudy (anonymous):

thank u tho

jhonyy9 (jhonyy9):

np was my pleasure - i think there is something wrong or i dont understand it right

OpenStudy (anonymous):

it's ok

jhonyy9 (jhonyy9):

@Help100 are you here again please ?

jhonyy9 (jhonyy9):

i have got it right now

jhonyy9 (jhonyy9):

look please on the diagram there is for x=1 y=4 yes ?

jhonyy9 (jhonyy9):

so than try f(x) -2 for x=1 how many will result for you ?

jhonyy9 (jhonyy9):

you know that f(x)=3*2^x so for x=1 how many will result f(1)= ?

jhonyy9 (jhonyy9):

f(1)=3*2^1 = ?

jhonyy9 (jhonyy9):

= 3*2=6 so than you see f(1) -2= 6-2=4 what mean that for x=1 result y=4 what there are exactly than you see on the diagram the right choice will be f(x)-2

OpenStudy (anonymous):

ok... thank u so much

jhonyy9 (jhonyy9):

i m happy that on the end i have got the right answer - choice - hope so much that you understand it now right sure - yes ?

jhonyy9 (jhonyy9):

ok. was my pleasure good luck bye and anytime than i m online bye